OpenStudy (anonymous):

Alex stores fertilizer for his farm in a rectangular box. The length of the box is 6 feet, width is 3 inches, and height is 9 inches. Each bag of fertilizer contains 0.225 cubic foot of fertilizer. How many bags of fertilizer will fill the box completely?

OpenStudy (anonymous):

I think it is five bags

OpenStudy (anonymous):

( length * width * height )/0.225 Refer to the attachment from Mathematica.

OpenStudy (anonymous):

6 feet = 72 inches 72*3*9 = 1944/0.225

OpenStudy (anonymous):

8640 ?? :o

OpenStudy (anonymous):

5 bags is what I calculated.

OpenStudy (anonymous):

lol yeah im thinking 5 bags too.

OpenStudy (anonymous):

because 12^3 = 1728

OpenStudy (anonymous):

and i divided 8640/1728 = 5

OpenStudy (anonymous):

Notice that in my pdf attachment that the numbers are alll integers or integer fractions. With this approach, the answers are always exact symbolic expressions.